Pricing Stock Issues in an IPO Zang Industries has hired the investment banking firm of Eric, Schwartz, & Mann (ESM) to helpplease make sure you got the right awenser i only got one chance to try
Answer #1

(a) Number of Shares Outstanding = 5 million, Market Value of Equity = $ 61 million, Price per Share = 61/5 = $ 12.2

Spread Charged = 5 % over and above the current price per share = 0.05 x 12.2 = $ 0.61

Therefore, Correct Offer Price = 12.2 + 0.61 = $ 12.81

(b) Number of New Shares Issued = 1.9 million, Offer Price = $ 12.81, Gross Proceeds = 12.81 x 1.9 = $ 24.339 million

Net Proceeds = Gross Proceeds - Underwriting Spread = 24.339 - 5 % of Gross Spread = 24.339 - 0.05 x 24.339 = $ 23.12205 million ~ $ 23.122 million